OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

bring it in y=mx+c form first them m is slope

mathslover (mathslover):

ok so .. first of all : 1) put x=0 , then y = ?

mathslover (mathslover):

2(0)+5y=-10 can u solve for y ?

OpenStudy (anonymous):

2x * times 0 = 0 so now we have 5y = -10 5y = -10 now we divide by 5 have y = 2????

mathslover (mathslover):

y = -10 / 5 =-2 check that again @SchoolFreak

OpenStudy (across):